A mining company receives samples of ore every 5 minutes in a Poissonian way, to which it orders under a PEPS type system, then the samples are analyzed. These analysis times behave under an exponential pattern with an average of 7 minutes. calculate: 1) the average number of samples waiting to be analyzed 2) the average time it takes for a sample in the laboratory
Answer #1

Given :

Arrival Rate () or Average number of Ore sample arrivals/hour = 60/5 = 12 samples/hour

Service rate, or Average samples serviced per hour = 60/7 = 8.57 samples/hour

1) Average number of samples waiting in line for analysis

= 2/(-)

= 122/8.57(8.57-12)

= -4.89 samples

2) Average time a sample spends in a system = Average time waiting in line + 1/

= /(-) + 1/

= 12/8.57(8.57-12) + 1/8.57

= -0.29 hours

Negative values indicate that the waiting line system is unstable and solution is not feasible.